### v2

# LLM Personas as a Substitute for Field Experiments in Method Benchmarking

Abstract
Field experiments (A/B tests) are often the most credible benchmark for methods (algorithms) in societal systems, but their cost and latency bottleneck rapid methodological progress. LLM-based persona simulation offers a cheap synthetic alternative, yet it is unclear whether replacing humans with personas preserves the benchmark interface that adaptive methods optimize against. We prove an if-and-only-if characterization: when (i) methods observe only the aggregate outcome (aggregate-only observation) and (ii) evaluation depends only on the submitted artifact and not on the method’s identity or provenance (method-blind evaluation), swapping humans for personas is just panel change from the method’s point of view, indistinguishable from changing the evaluation population (e.g., New York to Jakarta). Furthermore, we move from validity to usefulness: we define an information-theoretic discriminability of the induced aggregate channel and show that making persona benchmarking as decision-relevant as a field experiment is fundamentally a sample-size question, yielding explicit bounds on the number of independent persona evaluations required to reliably distinguish meaningfully different methods at a chosen resolution.
Machine Learning, ICML
1 Introduction
One of the recurring lessons from machine learning is that improvements in methods accelerate dramatically when we can iterate quickly against cheap and reliable evaluation benchmarks (Blum and Hardt, 2015; Zaharia et al., 2018; Liao et al., 2021; D’Amour et al., 2022; Bommasani et al., 2023; Miller, 2024; Abbas et al., 2025). The literature advances through an iterative process in which researchers propose a variant, test it against a benchmark, inspect failures, and refine. As a result, the limiting factor in progress is often the latency and reliability of the evaluation feedback loop rather than the availability of new ideas (Xin et al., 2018).
In many societal applications, such as pricing and matching policies, ad allocation, marketplace design, and behavioral interventions, the closest analogue to a benchmark is a field experiment (Harrison and List, 2004; Duflo, 2017), also known as A/B testing (Kohavi et al., 2020). Here, competing methods can be compared by deploying them through estimating their causal effects on agreed-upon outcome metrics. Repeating such experiments across settings allows researchers to assess not only average performance but also robustness and behavior under adaptation.
However, field experiments are costly to set up and slow to run (Diamond, 1986; Fisher, 1992; Paluck and Cialdini, 2014; Samek, 2019). Researchers typically must first collaborate with an organization, propose an intervention, negotiate approval across multiple stakeholders, and spend months implementing instrumentation and experimental infrastructure before any data are collected, and then wait to collect data until enough statistical power is gathered. Moreover, feasible experiments are often constrained by the organization’s interests, risk tolerance, and resource limitations, which restrict what methods can be tested and how frequently experiments can be repeated (Bandiera et al., 2011).
Recently, LLM-based persona simulation has shown potential to serve as a cheaper synthetic field laboratory (Toubia et al., 2025; Peng et al., 2025). In LLM-based persona simulation, a large language model is conditioned on explicit persona descriptions, such as survey responses that capture demographic attributes, preferences, roles, or beliefs. This is then used to generate individual-level evaluations or responses to proposed policies, content, or system outputs. These simulated micro-level responses can then be aggregated into a single evaluation signal. Because such simulations are inexpensive and fast to run at scale, they make it feasible to test and iterate on methods without field experimentation or A/B testing.
However, whether LLM-based persona simulation can indeed reliably substitute for field experimentation or A/B testing for methodology testing remains largely unknown. This uncertainty is exacerbated by recent studies that provide negative results in applications beyond methodology testing, particularly when we ask causal questions to support external validity claims (Gui and Toubia, 2023; Li et al., 2025; Peng et al., 2025; Gui and Kim, 2025). Manipulating prompts intended to represent a treatment can inadvertently shift other latent aspects of the scenario, inducing confounding.
Therefore, we face the following fundamental question:
When can LLM-based persona simulation serve as a drop-in substitute for a field experiment (or A/B test) as a benchmark for comparing methods?
In other words, under what protocol conditions is replacing humans with personas, from the benchmark’s perspective, equivalent to changing only the evaluation population (as when we change from the New York population to the Jakarta population)?
At first glance, this may sound like a question that can be empirically/experimentally validated: one might test persona benchmarking by checking whether persona scores correlate with human A/B test outcomes across a set of methods. However, such a correlation cannot validate or invalidate the ”drop-in substitute” claim. The claim is not about agreement of scores between personas and humans; it is about whether swapping the evaluator preserves the benchmark interface that methods optimize against. If personas simply represent a different evaluation population, then even in the best-case scenario, the induced score distribution can (and generally should) differ from the human one, just as the same method can score differently in New York than in Jakarta, so matching outcomes is neither necessary nor expected. For these reasons, the question we study is best treated as an identification question.
In this paper, we give the exact theoretical characterization of when this benchmark-level equivalence holds, by identifying two benchmark-hygiene conditions:
  • (i)
    aggregate-only observation (AO): Each method observes only the final aggregate score (and not the individual-level responses or individual identities),
  • (ii)
    method-blind evaluation (MB): For each method, the distribution of the returned score depends only on what was submitted and not on which training/optimization procedure produced.
We show that these two conditions are jointly necessary and sufficient for swapping human evaluation for persona evaluation to be equivalent to changing only the evaluation population, e.g., changing from the New York population to the Jakarta population. In other words, swapping humans for personas is indistinguishable by the method interface from an ordinary change of the evaluation population.
In addition to this identification result, we extend the identification discussion into a notion of usefulness: once (AO)+(MB) makes persona benchmarking identification-valid (i.e., “just panel change” on the method’s interface), the remaining question is whether the induced aggregate channel $Q_{\mathrm{pers}}(\cdot\mid w)$ is informative enough to distinguish and optimize meaningfully different methods. We formalize this via an information-based discriminability parameter (worst-case KL separation at a pre-specified resolution) and show it yields an explicit sample-complexity rule for how many independent persona evaluations (equivalently, what effective persona dataset/panel size) are required for reliable method comparison. In this sense, beyond enforcing (AO)+(MB), “persona quality” becomes a measurable budget question: is the persona panel large enough to resolve the improvements we care about?
The rest of the paper is organized as follows. Section 2 reviews related work. Section 3 formalizes the benchmarking setup and states the (AO) and (MB) conditions. Section 4 gives the main identification discussion, proving that if and only if (AO) and (MB) holds, persona vs. human evaluation is just panel change (JPC). Section 5 extends the identification discussion to a usefulness discussion that relates to sample complexity. The Appendix contains deferred proofs.
2 Related works
We discuss papers in the literature on benchmarks in AI in computer science and field experiment design that are related to the key contributions of our paper.
Aggregate-only observation (AO), a key condition for using persona simulation rather than field experiments for method benchmarking, has also appeared in the literature on benchmark gaming (Blum and Hardt, 2015; Hardt, 2017; Feldman et al., 2019; Biderman et al., 2024) as a baseline for preventing leaderboard gaming via adaptive overfitting. In econometrics, treatments of randomized experiments formalize the evaluation protocol as an assignment-and-aggregation procedure, clarifying what is identified from the realized (often aggregated) outcomes (Athey and Imbens, 2017).
Another key condition, method-blind evaluation (MB), has been experimentally motivated: people rate the same advice differently when it is labeled as AI rather than human/crowd-sourced (Bogert et al., 2022; Osborne and Bailey, 2025); many large-scale evaluation platforms implement partial blinding (e.g., anonymous pairwise comparisons) precisely to reduce provenance effects (Chiang et al., 2024). Dominguez-Olmedo et al. (2024) shows that survey-style elicitation of LLMs exhibits strong ordering and labeling effects that can dominate conclusions. The econometrics literature has also shown that outcomes can shift when participants condition on contextual/provenance cues, motivating blinding and stable scoring protocols (de Quidt et al., 2018; Levitt and List, 2011).
Related to quantifying a benchmark’s usefulness, Madaan et al. (2024) studied variance-focused analyses that quantify the extent to which benchmark scores fluctuate during evaluation. Heineman et al. (2025) discusses “signal vs. noise” perspectives and argue that benchmarks with higher signal-to-noise ratios are more reliable for model-selection decisions. Compared to these papers, we focus on proposing a benchmark-internal information measure (minimum KL separation) that directly yields sample-size scaling for reliable method comparison on the induced aggregate channel, clarifying how large a persona dataset is practically enough.
3 Setup: Benchmarking experimentation
3.1 Configurations and artifacts.
Let $\Theta$ denote the space of tunable configuration variables; this includes all controllable degrees of freedom that specify a system or a procedure (e.g., model weights, prompts/context, hyperparameters, decoding rules, tool policies, memory policies, data curation choices, or post-processing rules). A single candidate configuration is a choice $\theta\in\Theta$. A submitted artifact is what you hand to the benchmark, i.e., the externally visible object that the benchmark evaluates. We model this via an artifact map $g:\Theta\to\mathcal{W}$ and write
$w=g(\theta)\in\mathcal{W}.$
Possible choice of the artifact space $\mathcal{W}$ encompass: (i) a single output for a fixed input, (ii) an output distribution (stochastic method), (iii) a full interaction method mapping contexts to actions/outputs, (iv) a rollout distribution of an agent interacting with tools or environments.
Method as a configuration optimizer.
We model method benchmarking as a “submit–observe” loop. A method (algorithm) may be either non-adaptive (a single-shot submission, or a fixed distribution over submissions) or adaptive (updating submissions based on past benchmark feedback). Concretely, at each round $t=1,2,\dots,T$, the method chooses a configuration $\theta_{t}\in\Theta$ (equivalently an artifact $w_{t}:=g(\theta_{t})\in\mathcal{W}$), submits $w_{t}$ to an evaluator, and receives a feedback observation $o_{t}$ taking values in some observation space $\mathcal{O}$. Formally, a method (or method) $\mathcal{A}$ is modeled as a (possibly randomized) procedure that induces a decision kernel $\pi_{t}$ over configurations at each round $t$, such that
$\theta_{t}\sim\pi_{t}(\cdot\mid H_{t-1},S),$
where $H_{t-1}$ is the method’s observable history before round $t$, i.e.,
$H_{t-1}:=\{(\theta_{\tau},o_{\tau})\}_{\tau=1}^{t-1}$
and $S$ denotes any side information available before benchmarking begins (e.g., offline datasets, pretrained weights, logs, simulators).
Note that such a definition of an method contains:
  • Offline alignment (e.g., DPO (Rafailov et al., 2023), SFT (Wei et al., 2022) ) as the special case where $\pi_{t}$ does not depend on $H_{t-1}$ (or where $T=1$).
  • Online alignment (e.g., RLHF (Ouyang et al., 2022)) as the case where $\pi_{t}$ adapts to past feedback $o_{1:t-1}$.
  • Hyperparameter tuning / AutoML / architecture search (He et al., 2021; Ren et al., 2021) as black-box optimization over $\Theta$ using benchmark scores, where $\theta_{t}$ encodes a full training-and-deployment recipe, and $\pi_{t}$ implements a sequential search procedure.
  • Prompt and system configuration search (Pryzant et al., 2023; Kang and Yoganarasimhan, 2025) as the case where $\Theta$ indexes prompts, system messages, tool-use policies, retrieval and memory settings, and post-processing rules.
  • Data-centric training recipe search (Zha et al., 2025) as the case where $\Theta$ includes dataset construction and curation choices (filtering, reweighting, mixing, synthetic-data generation policies).
3.2 Evaluation as panel $\times$ instrument $\times$ aggregation
We model an evaluation as a two-stage process: many micro-level judgments are first produced, and these are then compressed into a single aggregate feedback signal that the method actually observes. Under this model, an evaluation setup (or simply an evaluator) is fully specified as
$(P,I,\Gamma,L),$
which we will treat as a primitive object throughout.
Panel ($P$).
A panel is a population of evaluators, either humans or LLM personas. Formally, let $\mathcal{P}$ denote the panel space and let $P$ be a distribution over $\mathcal{P}$. Each evaluation call draws a fresh panel of $L$ independent evaluators
$p_{1},\dots,p_{L}\overset{\text{i.i.d.}}{\sim}P.$
Micro-instrument ($I$).
Given an artifact $w\in\mathcal{W}$ and a panel member $p\in\mathcal{P}$, the micro-instrument produces an individual response. Formally, this is a conditional distribution
[…]
where the micro-response space $\mathcal{Z}$ can be a Likert score, a binary preference, or a short textual judgment. Each evaluator responds independently:
$Z_{\ell}\sim I(\cdot\mid w,p_{\ell}),\qquad\ell=1,\dots,L.$
Aggregation into observed feedback ($\Gamma$, $L$).
Finally, the benchmark aggregates the $L$ micro-responses into a single observable output. This is captured by a deterministic aggregation map
$\Gamma:\mathcal{Z}^{L}\to\mathcal{O},$ (2)
where $\mathcal{O}$ is the feedback observation space. Typical examples include the mean score, a majority vote, or a pass/fail indicator. Putting these pieces together, a single evaluation call on artifact $w$ returns the aggregate
$o\;=\;\Gamma(Z_{1},\dots,Z_{L})\in\mathcal{O}.$ (3)
What’s observed by the method.
Although evaluation involves panel members and micro-responses, the method never observes them. What it sees is only the induced distribution of the aggregate feedback $o\in\mathcal{O}$. The tuple $(P,I,\Gamma,L)$ therefore defines a Markov kernel on $\mathcal{O}$:
$Q_{P,I}(A\mid w)\;:=\;\mathbb{P}\big(\Gamma(Z_{1},\dots,Z_{L})\in A\,\big|\,w\big),$ (4)
$A\subseteq\mathcal{O}\ \text{measurable}.$
Intuitively, $Q_{P,I}(\cdot\mid w)$ is the distribution of the single observable feedback produced by the whole pipeline “sample panel $\to$ elicit micro-responses $\to$ aggregate” when the submitted artifact is $w$. All identification arguments in the sequel are necessarily about this reduced-form object.
3.3 Persona benchmark vs. human benchmark
Up to this point, we have described an evaluation setup abstractly as a tuple $(P,I,\Gamma,L)$. Now we instantiate this abstraction in the two cases we want to compare: evaluation by humans versus evaluation by LLM personas.
Human benchmark.
In a human benchmark, the panel distribution $P_{\mathrm{hum}}$ samples human evaluators, and the micro-instrument $I_{\mathrm{hum}}$ is the procedure that elicits a micro-response from a human (e.g., a rating, a preference, or a short written judgment). Together with the same aggregation map $\Gamma$ and panel size $L$, this induces an observable feedback kernel
$Q_{\mathrm{hum}}(\cdot\mid w).$
Persona benchmark.
In a persona benchmark, the panel distribution $P_{\mathrm{pers}}$ samples persona profiles (e.g., demographic or attitudinal descriptors), and the micro-instrument $I_{\mathrm{pers}}$ is implemented by an LLM judge conditioned on the sampled persona. Using the same aggregation $\Gamma$ and panel size $L$ yields a second observable feedback kernel
$Q_{\mathrm{pers}}(\cdot\mid w).$
What matters for the method.
Although these two pipelines differ internally (humans vs. personas; human judgments vs. LLM judgments), the method only interacts with each benchmark through the induced distribution of the aggregate feedback. In other words, for the method, the relevant objects are precisely the two reduced-form kernels $Q_{\mathrm{hum}}(\cdot\mid w)$ and $Q_{\mathrm{pers}}(\cdot\mid w)$.
3.4 Key assumptions
In asking the key question in this paper, “when can we treat persona evaluation as a clean benchmark interface for comparing methods?”, we need to discuss two “benchmark hygiene” conditions that clarify what information the method does (and does not) get access to, and whether the benchmark behaves like a well-defined environment independent of who is playing. They are not substantive modeling assumptions about humans or LLMs; they are assumptions about what information the benchmark reveals and how the evaluator behaves as an environment.
Assumption 1 (Aggregate-only observation (AO)).
At each round $t$, the method observes only the aggregate feedback $o_{t}\in\mathcal{O}$. It does not observe the micro-level tuple $(p_{1},\dots,p_{L},Z_{1},\dots,Z_{L})$, any panel identifiers, or any additional side-channel information beyond $o_{t}$.
Intuitively, (AO) says the method sees exactly what a standard leaderboard would show: one score (or label) per submission. This prevents method’s gaming behavior that relies on recognizing individual panelists/personas or exploiting micro-level structure that would be invisible in the intended benchmark interface.
Before discussing the next assumption, we define a probability measure $\mathbb{P}^{\mathcal{A}}$: for a fixed benchmark implementation, running an method $\mathcal{A}$ induces a probability measure $\mathbb{P}^{\mathcal{A}}$ over the interaction transcript $(w_{1},o_{1},w_{2},o_{2},\dots)$. Under (AO), we can define the method’s interaction transcript after t1t-1 rounds as
$\widetilde{H}_{t-1}:=\{(w_{\tau},o_{\tau})\}_{\tau=1}^{t-1},$
and its information before choosing $w_{t}$ is the $\sigma$-field
$\mathcal{I}_{t-1}:=\sigma(S,\mathsf{R},\widetilde{H}_{t-1}).$
Assumption 2 (Method-blind evaluation (MB)).
There exists a Markov kernel $Q(\cdot\mid w)$ on $\mathcal{O}$ such that for every method $\mathcal{A}$, every round $t$, and every measurable $A\subseteq\mathcal{O}$,
$\mathbb{P}^{\mathcal{A}}\!\big(o_{t}\in A\,\big|\,\mathcal{I}_{t-1},\,w_{t}\big)=Q(A\mid w_{t})\quad\text{a.s.}$
Intuitively, (MB) is the minimal condition for calling this evaluation setup a benchmark environment at all: the evaluator should not care about the identity of the training procedure or other metadata, and care only about what was submitted (the artifact). In other words, the benchmark interaction is fully summarized by the reduced-form kernel $Q(\cdot\mid w)$, which is fixed across methods.
4 Identification: When Is Persona Benchmarking “Just Panel Change”?
From the method’s point of view, each benchmark is a black box: it takes an artifact $w\in\mathcal{W}$ and returns a random aggregate feedback value $o\in\mathcal{O}$. All the internal structure (panel draws, micro-judgments, and aggregation) has already been compressed into the reduced-form kernels
$Q_{\mathrm{pers}}(\cdot\mid w)\quad\text{and}\quad Q_{\mathrm{hum}}(\cdot\mid w),$
defined in (4). In this section, we utilize this intuition to answer to the following question:
When is swapping human evaluation for persona evaluation, as seen through the method’s interface, nothing more than changing the panel $P$?
We first formalize what it means to “only change the evaluation population” in the panel–instrument–aggregation model (a literal panel change), and then define the corresponding interface-level notion that is relevant for adaptive benchmarking methods (just panel change, JPC). We then show that for human vs. persona benchmarking, JPC holds if and only if two benchmark-hygiene conditions hold: aggregate-only observation (AO) and method-blind evaluation (MB).
This yields an identification result: under (AO)+(MB), swapping humans for personas is indistinguishable on the method interface from an ordinary change of evaluation population (even though the internal micro-instrument may differ); conversely, if either condition fails, the swap can change the interface in ways that go beyond “panel change.”
4.1 Literal panel change iff just panel change (JPC)
Definition 1 (Literal panel change).
Fix a panel space $\mathcal{P}$, micro-response space $\mathcal{Z}$, aggregation map $\Gamma:\mathcal{Z}^{L}\to\mathcal{O}$ and panel size $L\in\mathbb{N}$. Let $I(\cdot\mid w,p)$ be a micro-instrument on $\mathcal{Z}$. For two panel distributions $P,P^{\prime}$ on $\mathcal{P}$, define two benchmarks
$B:=(P,I,\Gamma,L),\qquad B^{\prime}:=(P^{\prime},I,\Gamma,L).$
We say $B^{\prime}$ is obtained from $B$ by a literal panel change if the only difference between them is that $P$ is replaced by $P^{\prime}$ (i.e., $I,\Gamma,L$ are identical).
Definition 1 captures the classical “same survey, different respondents” intuition: two benchmarks has the same instrument (the question/rubric and how responses are generated), the same aggregation rule, and the same sample size, but we sample from a different population (e.g., one from New York and one from Jakarta). That is, the protocol is unchanged except for the distribution. over who evaluates.
Swapping humans for personas is not literally a panel change in this narrow sense because the micro-instrument is implemented differently (humans vs. an LLM judge conditioned on a persona). This motivates an interface-level notion that asks whether the swap behaves like a panel change to any method that can adapt to benchmark feedback; the following Just panel change (JPC) definition serves the role.
Definition 2 (Just panel change (JPC)).
Consider two benchmarks $B$ and $B^{\prime}$ on the same artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. We say swapping $B$ for $B^{\prime}$ is just panel change (JPC) on the method interface if there exist Markov kernels $Q:\mathcal{W}\mapsto\Delta\mathcal{O}$ and $Q^{\prime}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that, for every method $\mathcal{A}$ and every horizon $T$, the observable transcript laws factorize as
$\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T})$ $=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q(do_{t}\mid w_{t}),$ (5)
$\mathbb{P}^{\mathcal{A}}_{B^{\prime}}(dw_{1:T},do_{1:T})$ $=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q^{\prime}(do_{t}\mid w_{t}).$ (6)
In words: for every method, the swap preserves the observation/information structure and differs only in their kernels $Q_{\mathrm{hum}}$ and $Q_{\mathrm{pers}}$.
Definition 2 is intentionally method-centric: it quantifies over all (possibly adaptive) methods and asks whether, from their perspective, the benchmark is an oracle channel that depends only on the current submission $w_{t}$, with no additional side channels or method-dependent behavior. Crucially, JPC does not require $Q=Q^{\prime}$ (scores need not match across humans and personas); it only requires that the form of the interaction is preserved and that the swap can be summarized entirely by replacing one artifact-to-score kernel by another. This is precisely the sense in which swapping evaluators should resemble a “panel change” rather than a change in the rules of the game; the following Lemma 4.1 and 4.2 formalize the mathematical equivalence.
Lemma 4.1 (Literal panel change $\Rightarrow$ JPC).
Assume Aggregate-only observation (AO), i.e., the method observes only $o_{t}\in\mathcal{O}$ each round. Let $B=(P,I,\Gamma,L)$ and $B^{\prime}=(P^{\prime},I,\Gamma,L)$ differ by a literal panel change in the sense of Definition 1. Define the reduced-form kernels
$Q_{P,I}(A\mid w)$ $:=\mathbb{P}\big(\Gamma(Z_{1},\dots,Z_{L})\in A\mid w\big),$
$Q_{P^{\prime},I}(A\mid w)$ $:=\mathbb{P}\big(\Gamma(Z^{\prime}_{1},\dots,Z^{\prime}_{L})\in A\mid w\big),$
as in (4). Then the swap $B\leftrightarrow B^{\prime}$ is JPC in the sense of Definition 2, with $Q:=Q_{P,I}$ and $Q^{\prime}:=Q_{P^{\prime},I}$.
Lemma 4.2 (JPC $\Rightarrow$ observational equivalence to a literal panel change).
Suppose two benchmarks $B$ and $B^{\prime}$ satisfy JPC in the sense of Definition 2, i.e., there exist Markov kernels $Q,Q^{\prime}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$ and horizon $T$,
$\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T})$ $=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q(do_{t}\mid w_{t}),$
$\mathbb{P}^{\mathcal{A}}_{B^{\prime}}(dw_{1:T},do_{1:T})$ $=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q^{\prime}(do_{t}\mid w_{t}).$
Then there exists a pair of benchmarks $\overline{B}=(\overline{P},\overline{I},\overline{\Gamma},\overline{L})$ and $\overline{B}^{\prime}=(\overline{P}^{\prime},\overline{I},\overline{\Gamma},\overline{L})$ that differ by a literal panel change (i.e., only $\overline{P}$ is replaced by $\overline{P}^{\prime}$), such that for every method $\mathcal{A}$ and horizon $T$,
$\mathbb{P}^{\mathcal{A}}_{\overline{B}}=\mathbb{P}^{\mathcal{A}}_{B},\qquad\mathbb{P}^{\mathcal{A}}_{\overline{B}^{\prime}}=\mathbb{P}^{\mathcal{A}}_{B^{\prime}}.$
In particular, JPC is exactly the statement that the swap is indistinguishable on the method interface from a literal panel change.
Lemma 4.1 says that, if you literally change only the panel distribution $P$ in the panel–instrument–aggregation model, then the observable transcript laws change only by replacing the reduced-form kernel $Q_{P,I}$ by $Q_{P^{\prime},I}$, i.e., JPC holds. Lemma 4.2 says the converse at the interface level: any JPC swap can be realized exactly as a literal panel change in a (possibly abstract) benchmark representation. Thus Definition 2 is mathematically equivalent to “literal panel change” as an interface-level notion (i.e., up to equality of transcript laws for every adaptive method).
The next subsection provides a protocol-level characterization of exactly when that interface property holds.
4.2 Just panel change (JPC) iff (AO)+(MB)
Lemma 4.3 ((JPC) $\iff$ (AO)+(MB)).
Let $B_{\mathrm{hum}}$ and $B_{\mathrm{pers}}$ denote the human and persona benchmarking protocols, with common artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. The following are equivalent:
  1. 1.
    (JPC). The human $\leftrightarrow$ persona swap is just panel change (JPC) on the method interface (Definition 2).
  2. 2.
    (AO)+(MB). The protocol satisfies aggregate-only observation (AO) (Assumption 1) and each benchmark is method-blind (MB) (Assumption 2).
At a technical level, JPC is defined by a factorization of the transcript laws through an artifact-only kernel. This makes the main theorem appear almost tautological: if we assume conditions that guarantee exactly such a factorization, we recover JPC. The point of stating the equivalence is not to rebrand a definition, but to (i) express JPC in terms of two concrete and auditable benchmark-hygiene requirements (what the benchmark reveals, and whether it is provenance-blind), and (ii) cleanly separate the two ways persona benchmarking can fail as a substitute for field experiments: leakage of micro-level information (violating AO) and method/provenance dependence (violating MB).
Lemma 4.3 also highlights two distinct failure modes:
If (MB) fails, the benchmark is not a well-defined oracle environment.
When (MB) fails, there is no single kernel $Q(\cdot\mid w)$ that governs the returned score across methods. Equivalently, the evaluator’s behavior depends on provenance/identity or on interaction history in a way that is not summarized by the submitted artifact. In such a case, the benchmark cannot be treated as an artifact-only oracle channel, so “panel change” is not an identified description of the swap. See Appendix B.2 for when MB can be justified.
If (AO) fails, the method interface changes (even if aggregate scores look the same).
If the protocol reveals micro-level information (panel identities, raw votes, ordering, etc.), then two evaluator implementations can induce the same aggregate kernel on $\mathcal{O}$ but still be distinguishable and exploitable by an adaptive method. Hence the observable interaction is not characterized by the aggregate channel alone. See Appendix B.1 for when AO fails.
5 Beyond validity: When does a persona panel constitute an useful benchmark?
As we discussed in section 4, (AO)+(MB) characterizes when and only when persona benchmarking is a valid benchmark, satisfying (JPC). But validity, which is an identification argument, is not necessarily equivalent to usefulness. In this section, we formalize when and when not a persona-based LLM simulation can be a perfectly valid, but much less useful, benchmark compared to field experiments. We show that 1) dataset size is what really matters for a persona-based benchmark, or any benchmark, to be considered a useful benchmark, and 2) how to find such a required dataset size empirically.
As a starting point, recall that, under (AO)+(MB), the method interacts with the persona benchmark only through the reduced-form kernel
$Q_{\mathrm{pers}}(\cdot\mid w)\quad\text{on}\quad\mathcal{O},$
which returns an aggregate feedback draw $o\sim Q_{\mathrm{pers}}(\cdot\mid w)$ for each submitted artifact $w$. Thus, we arrive at an important observation:
Whether a persona dataset will constitute an useful benchmark is a question about how informative the induced channel $Q_{\mathrm{pers}}$ is for distinguishing and optimizing methods.
Throughout this section, we will utilize this idea to derive the minimum persona dataset size for benchmarking.
5.1 Usefulness and “mountain-fog” analogy
Denote
$\mathcal{W}_{0}\subseteq\mathcal{W}$
as the region of interest for the artifacts (e.g., a tunable region of artifacts around a baseline artifact). Assume that $\mathcal{O}=\mathbb{R}$ (a scalar score), define the benchmark landscape and benchmark noise by
$\mu_{\mathrm{pers}}(w):=\mathbb{E}[\,o\mid w\,],\quad\sigma_{\mathrm{pers}}^{2}(w):=\mathrm{Var}(o\mid w),$ (7)
where $o\sim Q_{\mathrm{pers}}(\cdot\mid w).$
A useful analogy to describe the usefulness of a benchmark is the “mountain-fog” metaphor. The (unknown) benchmark landscape $w\mapsto\mu_{\mathrm{pers}}(w)$ is the mountain: it assigns to each artifact the expected aggregate score returned by the benchmark. The benchmark noise scale $w\mapsto\sigma_{\mathrm{pers}}(w)$ is the fog. This metaphor separates two different reasons a benchmark may not be useful. A benchmark can be a flat mountain, meaning that $\mu_{\mathrm{pers}}(w)$ changes only slightly across meaningfully different artifacts; or it can be noisy (thick fog), meaning that $\sigma_{\mathrm{pers}}(w)$ is large relative to the score differences the method is trying to detect.
5.2 Formalization of usefulness: discriminability
We formalize the “mountain-fog” idea by introducing a concept we call discriminability. let $D_{\text{KL}}(\cdot,\cdot)$ be Kullback–Leibler divergence on $\mathcal{O}$. Also, suppose that we can define a metric $d_{\mathcal{W}}$ on $\mathcal{W}_{0}$.
We define the discriminability of a benchmark by
$\kappa_{Q}:=\;$
$\;\inf_{\begin{subarray}{c}w,w^{\prime}\in\mathcal{W}_{0}\\ d_{\mathcal{W}}(w,w^{\prime})\geq r\end{subarray}}D_{\text{KL}}\!\left(Q_{\mathrm{pers}}(\cdot\mid w),\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})\right).$ (8)
where $r>0$ is a resolution parameter that formalizes “minimal meaningful change” in artifacts, which is often pre-specified ex ante. The pair $(d_{\mathcal{W}},r)$ should be read as: “we only require the benchmark to separate artifacts that differ by at least $r$ under $d_{\mathcal{W}}$.”
Choosing a smaller $r$ is a stricter requirement (it asks the benchmark to resolve finer changes), and it can only make discriminability harder (the infimum ranges over a larger set, so $\kappa_{Q}$ can only decrease). Thus $r$ should reflect the smallest change that is substantively meaningful for method development, not the smallest change that can be expressed syntactically. In Section 5.4, we discuss the guidelines of figuring out $d_{\mathcal{W}}$ and $r$. If $\kappa_{Q}$ is near zero, it means that distinct artifacts in $\mathcal{W}_{0}$ that are essentially indistinguishable through the benchmark interface, and therefore, the benchmark provides little usable feedback for methods.
In practice, the KL divergence is challenging to estimate empirically. The following Lemma 5.1 provides a nice simplification under the canonical assumption of homoscedastic Gaussian reduced-form kernels, resolving this challenge.
Lemma 5.1.
Assume $\mathcal{O}=\mathbb{R}$ and that a benchmark induces a Gaussian homoscedastic reduced-form kernel:
$Q_{\mathrm{pers}}(\cdot\mid w)=\mathcal{N}\!\big(\mu_{\mathrm{pers}}(w),\,\sigma^{2}\big),\quad\sigma>0$ (9)
Then for any $w,w^{\prime}\in\mathcal{W}$,
$D_{\mathrm{KL}}\!\big(Q_{\mathrm{pers}}(\cdot\mid w)\,\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})\big)$
$=\;\frac{\big(\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})\big)^{2}}{2\sigma^{2}}=\,\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}.$ (10)
Moreover, for a dataset with population size $L$,
$D_{\mathrm{KL}}\!\big(Q_{\mathrm{pers}}(\cdot\mid w)^{\otimes L}\,\big\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})^{\otimes L}\big)=L\,\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}.$
Note that the right hand side of Equation (10) is a quantity we often call the pairwise signal-to-noise (SNR), which is defined as
$\mathrm{SNR}(w,w^{\prime})\;:=\;\frac{\Delta(w,w^{\prime})^{2}}{\sigma_{\mathrm{pers}}^{2}(w)+\sigma_{\mathrm{pers}}^{2}(w^{\prime})}=\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}$
Therefore, by combining Equation (8) and (10), $\kappa_{Q}$ becomes
$\kappa_{Q}=\inf_{\begin{subarray}{c}w,w^{\prime}\in\mathcal{W}_{0}\\ d_{\mathcal{W}}(w,w^{\prime})\geq r\end{subarray}}\mathrm{SNR}(w,w^{\prime})$ (11)
Note that Equation (11) empirically estimable: the Pairwise signal-to noise ratio $\operatorname{SNR}\left(w,w^{\prime}\right)$ depends only on the first two moments of this kernel $Q$, namely the mean $\mu(w)=\mathbb{E}[o\mid w]$ and variance $\sigma^{2}(w)=\operatorname{Var}(o\mid w)$. Both quantities are identifiable from repeated independent calls to the benchmark at a fixed artifact $w$, by re-running the same evaluation protocol with i.i.d. samples $o\sim Q(\cdot\mid w)$.
5.3 Sample complexity.
The discriminability $\kappa_{Q}$ defined in (8) has a direct operational interpretation: it is the per-sample information available to distinguish two artifacts that differ by at least $r$ in the benchmark interface. Consequently, the number of persona samples required for a reliable pairwise comparison scales inversely with $\kappa_{Q}$.
Lemma 5.2 (Pairwise comparison sample complexity from discriminability).
Assume Gaussian model (9). Fix $r>0$ and let $\kappa_{Q}$ be as in (8). For any $w,w^{\prime}\in\mathcal{W}_{0}$ with $d_{\mathcal{W}}(w,w^{\prime})\geq r$ and $\Delta(w,w^{\prime}):=\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})>0$, evaluate each artifact $L$ times with i.i.d. fresh panel draws (equivalently, i.i.d. draws $o\sim Q_{\mathrm{pers}}(\cdot\mid w)$) and compare empirical means. Then
$\mathbb{P}\!\left(\widehat{\mu}_{n}(w)\leq\widehat{\mu}_{n}(w^{\prime})\right)\leq\exp\!\left(-\frac{L}{2}\frac{\,\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}\right)$
$\leq\exp\!\left(-\frac{L}{2}\,\kappa_{Q}\right).$
In particular, in probability at most $\delta$, it suffices that the dataset size $L$ satisfies
$L\;\geq\;\frac{2}{\kappa_{Q}}\log\frac{1}{\delta}.$
Lemma 5.2 states that $\kappa_{Q}$ can be understood as the “per-data” information available to distinguish two meaningfully different artifacts (separated by at least $r$). When $\kappa_{Q}$ is small, because the benchmark is flat (small $\Delta$) or noisy (large $\sigma^{2}$), then any method requires many independent data samples to reliably decide whether one artifact is better than another. That is, this scaling clarifies the effective size of dataset for benchmarking.
5.4 Choice of $r$ and $d_{\mathcal{W}}$
The definition of discriminability in (11) depends on two user-specified design choices: a metric $d_{\mathcal{W}}$ on the artifact space $\mathcal{W}_{0}$ and a resolution threshold $r>0$. These are generally method- and task-specific design parameters: different method families explore different degrees of freedom in $\mathcal{W}$ and therefore induce different natural notions of distance and resolution. Operationally, $d_{\mathcal{W}}$ and $r$ determine which pairs of artifacts the benchmark is required to reliably distinguish, and therefore they determine the relevant sample complexity via Lemma 5.2.
Below are practical guidelines for selecting them in a way that is both interpretable and robust.
Tie $d_{\mathcal{W}}$ to the developer’s degrees of freedom.
A good default is to define $d_{\mathcal{W}}$ via the natural parameterization that methods actually tune. If artifacts are produced by knobs $\theta\in\Theta$ through $w(\theta)$, and there is a natural distance $d_{\Theta}$ on $\Theta$, one can induce a pseudo-metric on $\mathcal{W}$ by
$d_{\mathcal{W}}(w(\theta),w(\theta^{\prime})):=d_{\Theta}(\theta,\theta^{\prime}).$
This makes $r$ interpretable as a step size in the space the method explores. Examples include the scaled Euclidean distance on continuous hyperparameters or the edit distance on a structured prompt template.
Choose $r$ as a minimal meaningful iteration unit.
In most benchmarking use cases, there is a natural notion of the smallest “iteration” a developer expects to be worth distinguishing. The guiding principle is that $r$ should be large enough that changes below $r$ are not worth spending benchmark budget on, but small enough that improvements developers actually seek fall above $r$
  • Prompt/instruction tuning: $r$ can be “one allowed edit” under a pre-specified edit set (add/remove one constraint, add one example, modify one rubric item). Under token-level edit distance, this corresponds to a small fixed number of edits.
  • Hyperparameter tuning: choose a scaled metric so that a standard “one-step” change has size $\approx 1$, then set $r=1$. For instance, scale each coordinate by a typical tuning increment.
  • Model or policy variants: set $r$ to the smallest recipe change you would treat as a distinct method (e.g., one additional fine-tuning epoch, one dataset mixture adjustment above a threshold, a decoding rule change).
In short, $d_{\mathcal{W}}$ should encode meaningful artifact differences (preferably behavioral and invariant to cosmetic changes), while $r$ should encode the smallest change that developers intend to reliably resolve. With these choices fixed, $\kappa_{Q}$ becomes an operationally estimable quantity, and Lemma 5.2 translates it directly into the required persona data size for stable method comparison.
5.5 Experiments
As a proof-of-concept, Appendix C demonstrates how our discriminability-based analysis can be used to calibrate the persona dataset size needed for reliable method comparison. Concretely, we apply the procedure to the persona-simulation ad benchmark used to evaluate TextBO, a prompt-optimization based self-improving AI system (Kang and Yoganarasimhan, 2025), and compute the implied number of independent persona evaluations required to distinguish one-step prompt improvements at a chosen confidence level.
6 Conclusion
We characterized when LLM-persona panels can substitute for human field experiments as a benchmark interface for method development. Our main result shows that persona vs. human evaluation is just panel change from the method’s perspective if and only if two benchmark-hygiene conditions hold: (i) aggregate-only observation (AO) and (ii) method-blind evaluation (MB). When either fails, the benchmark can leak exploitable information or depend on provenance, breaking the interface-level equivalence.
We also separated validity from usefulness. A valid persona benchmark may still be less informative than field experimentation if the induced aggregate channel is too flat or noisy. Our discriminability $\kappa_{Q}$ (worst-case KL separation at resolution $r$) yields the corresponding budget scaling for reliable comparisons, on the order of $\kappa_{Q}^{-1}\log(1/\delta)$. Thus, beyond enforcing (AO)+(MB), the practical requirement is a sufficient persona dataset.
Appendix A Deferred theoretical discussions
A.1 Proof of Lemma 4.1 and 4.2 in Section 4
Proof of Lemma 4.1.
Fix any method $\mathcal{A}$ with submission kernels $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO), and fix a round $t$. Under benchmark $B=(P,I,\Gamma,L)$, conditional on the submitted artifact $w_{t}$, the benchmark generates
$p_{t,1},\dots,p_{t,L}\overset{i.i.d.}{\sim}P,\qquad Z_{t,\ell}\sim I(\cdot\mid w_{t},p_{t,\ell})\ \text{independently over }\ell,\qquad o_{t}=\Gamma(Z_{t,1},\dots,Z_{t,L}).$
By construction, given $w_{t}$ this sampling uses only fresh benchmark randomness (fresh panel draw and micro-responses) and therefore does not depend on $\mathcal{I}_{t-1}$ nor on the identity of $\mathcal{A}$. Hence for every measurable $A\subseteq\mathcal{O}$,
$\mathbb{P}_{B}^{\mathcal{A}}(o_{t}\in A\mid\mathcal{I}_{t-1},w_{t})=Q_{P,I}(A\mid w_{t})\quad\text{a.s.}$
This is exactly the (MB)-type conditional independence statement with kernel $Q_{P,I}$. Combining this with the fact that $w_{t}\sim\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO), the standard sequential composition / chain rule for Markov kernels yields, for every horizon $T$,
$\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{P,I}(do_{t}\mid w_{t}).$
The same argument for $B^{\prime}=(P^{\prime},I,\Gamma,L)$ gives
$\mathbb{P}^{\mathcal{A}}_{B^{\prime}}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{P^{\prime},I}(do_{t}\mid w_{t}).$
This is precisely Definition 2 with $Q:=Q_{P,I}$ and $Q^{\prime}:=Q_{P^{\prime},I}$. ∎
Proof of Lemma 4.2.
We construct an auxiliary panel–instrument–aggregation representation that reproduces the same transcript laws. This construction is not meant to mirror the internal structure of the original human/persona protocols. Let
$\overline{\mathcal{P}}$ $:=\{0,1\},\overline{\mathcal{Z}}:=\mathcal{O},\ \overline{L}:=1,$
$\overline{\Gamma}$ $:\overline{\mathcal{Z}}\to\mathcal{O}\ \text{be the identity map }\overline{\Gamma}(o)=o.$
Define a single micro-instrument $\overline{I}$ on $\overline{\mathcal{Z}}=\mathcal{O}$ by
$\overline{I}(\cdot\mid w,0):=Q(\cdot\mid w),\qquad\overline{I}(\cdot\mid w,1):=Q^{\prime}(\cdot\mid w).$
Now define the two panel distributions
$\overline{P}:=\delta_{0},\qquad\overline{P}^{\prime}:=\delta_{1}.$
Then $\overline{B}:=(\overline{P},\overline{I},\overline{\Gamma},\overline{L})$ and $\overline{B}^{\prime}:=(\overline{P}^{\prime},\overline{I},\overline{\Gamma},\overline{L})$ differ only in the panel distribution ($\delta_{0}$ versus $\delta_{1}$), hence are a literal panel change.
Moreover, under $\overline{B}$, each benchmark call on $w$ samples $p=0$ a.s., then outputs $o\sim\overline{I}(\cdot\mid w,0)=Q(\cdot\mid w)$. Thus the reduced-form kernel of $\overline{B}$ is exactly $Q$. Similarly, the reduced-form kernel of $\overline{B}^{\prime}$ is exactly $Q^{\prime}$. Therefore, for every method $\mathcal{A}$ and horizon $T$, the transcript laws under $\overline{B}$ and $\overline{B}^{\prime}$ satisfy the same factorizations as in the displayed JPC equations, which implies
$\mathbb{P}^{\mathcal{A}}_{\overline{B}}(dw_{1:T},do_{1:T})$
$=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q(do_{t}\mid w_{t})=\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T}),$
and likewise $\mathbb{P}^{\mathcal{A}}_{\overline{B}^{\prime}}=\mathbb{P}^{\mathcal{A}}_{B^{\prime}}$. ∎
A.2 Proof of Lemma 4.3 in Section 4
Proof of Lemma 4.3.
We prove $(2)\Rightarrow(1)$ and $(1)\Rightarrow(2)$.
$(2)\Rightarrow(1)$: (AO)+(MB) imply JPC. Assume (AO) holds and each benchmark is method-blind (MB) in the sense of Assumptions 12. Apply Lemma A.1 to the human benchmark $B_{\mathrm{hum}}$ to obtain a kernel $Q_{\mathrm{hum}}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$ and horizon $T$,
$\mathbb{P}^{\mathcal{A}}_{B_{\mathrm{hum}}}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{\mathrm{hum}}(do_{t}\mid w_{t}).$
Likewise apply Lemma A.1 to $B_{\mathrm{pers}}$ to obtain $Q_{\mathrm{pers}}$ with
$\mathbb{P}^{\mathcal{A}}_{B_{\mathrm{pers}}}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{\mathrm{pers}}(do_{t}\mid w_{t}).$
Setting $Q:=Q_{\mathrm{hum}}$ and $Q^{\prime}:=Q_{\mathrm{pers}}$ verifies Definition 2. Hence the swap is JPC on the method interface.
$(1)\Rightarrow(2)$: JPC implies (AO)+(MB). Assume the swap is JPC in the sense of Definition 2. By that definition, the transcript laws of both benchmarks factorize for every method $\mathcal{A}$ and every horizon $T$ through some kernels $Q$ and $Q^{\prime}$.
First, (AO) is the observation/information structure assumed in Definition 2 (the method’s interaction is summarized by the aggregate-history $\mathcal{I}_{t-1}$ and transcript $(w_{t},o_{t})$). Second, applying Lemma A.2 to $B_{\mathrm{hum}}$ and $B_{\mathrm{pers}}$ shows that each benchmark satisfies (MB) with kernels $Q_{\mathrm{hum}}=Q$ and $Q_{\mathrm{pers}}=Q^{\prime}$ respectively.
Therefore, (AO)+(MB) hold. ∎
Lemma A.1 (Transcript factorization under (AO)+(MB)).
Fix a benchmark $B$ with artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. Assume (AO) and (MB) hold for $B$, i.e., there exists a Markov kernel $Q_{B}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$, every round $t$, and every measurable $A\subseteq\mathcal{O}$,
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid w_{t},H_{t-1})=Q_{B}(A\mid w_{t})\quad\text{a.s.}$ (12)
Then for every method $\mathcal{A}$ (with submission kernels $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO)) and every horizon $T$,
$\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{B}(do_{t}\mid w_{t}),$ (13)
in the usual sequential composition-of-kernels sense.
Proof of Lemma A.1.
Fix an method $\mathcal{A}$ and a round $t$.
Step 1. Since $H_{t-1}\subseteq\mathcal{I}_{t-1}$ and $Q_{B}(A\mid w_{t})$ is measurable w.r.t. $\sigma(w_{t})$, we have by the tower property:
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid\mathcal{I}_{t-1},w_{t})$ $=\mathbb{E}^{\mathcal{A}}_{B}\!\left[\mathbf{1}\{o_{t}\in A\}\mid\mathcal{I}_{t-1},w_{t}\right]$
$=\mathbb{E}^{\mathcal{A}}_{B}\!\left[\mathbb{E}^{\mathcal{A}}_{B}\!\left[\mathbf{1}\{o_{t}\in A\}\mid H_{t-1},w_{t}\right]\;\middle|\;\mathcal{I}_{t-1},w_{t}\right]$
$=\mathbb{E}^{\mathcal{A}}_{B}\!\left[Q_{B}(A\mid w_{t})\mid\mathcal{I}_{t-1},w_{t}\right]$
$=Q_{B}(A\mid w_{t}),$
a.s. Hence, conditional on $(\mathcal{I}_{t-1},w_{t})$, the benchmark draw $o_{t}$ is distributed as $Q_{B}(\cdot\mid w_{t})$.
Step 2. Under (AO), the method selects $w_{t}$ using the kernel $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$. Combining with Step 1, conditional on $\mathcal{I}_{t-1}$ we can generate $(w_{t},o_{t})$ by:
$w_{t}\sim\pi_{t}(\cdot\mid\mathcal{I}_{t-1}),\qquad o_{t}\sim Q_{B}(\cdot\mid w_{t}).$
Step 3. Applying the standard chain rule / iterated conditioning for Markov kernels yields the full-horizon factorization (13):
$\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{B}(do_{t}\mid w_{t}).$
Lemma A.2 (Factorization implies (MB)).
Fix a benchmark $B$. Suppose that under (AO), there exists a Markov kernel $Q_{B}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$ and every horizon $T$,
$\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{B}(do_{t}\mid w_{t}).$
Then $B$ satisfies (MB) in the sense of Assumption 2; namely, for every method $\mathcal{A}$, every $t$, and measurable $A\subseteq\mathcal{O}$,
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid w_{t},H_{t-1})=Q_{B}(A\mid w_{t})\quad\text{a.s.}$
Proof of Lemma A.2.
Fix $\mathcal{A},t$ and a measurable $A\subseteq\mathcal{O}$.
Under the assumed factorization, conditional on $\mathcal{I}_{t-1}$ and $w_{t}$, the draw $o_{t}$ is generated by $Q_{B}(\cdot\mid w_{t})$, so
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid\mathcal{I}_{t-1},w_{t})=Q_{B}(A\mid w_{t})\quad\text{a.s.}$
Now apply the tower property conditioning down to $(H_{t-1},w_{t})$:
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid H_{t-1},w_{t})$ $=\mathbb{E}^{\mathcal{A}}_{B}\!\left[\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid\mathcal{I}_{t-1},w_{t})\;\middle|\;H_{t-1},w_{t}\right]$
$=\mathbb{E}^{\mathcal{A}}_{B}\!\left[Q_{B}(A\mid w_{t})\mid H_{t-1},w_{t}\right]$
$=Q_{B}(A\mid w_{t}),$
since $Q_{B}(A\mid w_{t})$ is $\sigma(w_{t})$-measurable. This is exactly (MB). ∎
A.3 Proofs in Section 5
Proof of Lemma 5.1..
Let $P_{x}$ denote the density of $\mathcal{N}(x,\sigma^{2})$:
$p_{x}(o)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\!\left(-\frac{(o-x)^{2}}{2\sigma^{2}}\right).$
By definition,
$d(x,y)=\mathbb{E}_{O\sim\mathcal{N}(x,\sigma^{2})}\!\left[\log\frac{p_{x}(O)}{p_{y}(O)}\right].$
Compute the log-likelihood ratio:
$\log\frac{p_{x}(O)}{p_{y}(O)}$ $=-\frac{(O-x)^{2}}{2\sigma^{2}}+\frac{(O-y)^{2}}{2\sigma^{2}}$
$=\frac{(O-y)^{2}-(O-x)^{2}}{2\sigma^{2}}$
$=\frac{(x-y)\big(2O-x-y\big)}{2\sigma^{2}}.$
Taking expectation under $O\sim\mathcal{N}(x,\sigma^{2})$ yields
$d(x,y)$ $=\frac{x-y}{2\sigma^{2}}\,\mathbb{E}\big[2O-x-y\big]$
$=\frac{x-y}{2\sigma^{2}}\,(2x-x-y)=\frac{(x-y)^{2}}{2\sigma^{2}},$
Now note that $O_{1:L}$ are i.i.d. from $\mathcal{N}(x,\sigma^{2})$, the joint density is $\prod_{i=1}^{L}p_{x}(O_{i})$, so the joint log-likelihood ratio satisfies
$\log\frac{\prod_{i=1}^{L}p_{x}(O_{i})}{\prod_{i=1}^{L}p_{y}(O_{i})}=\sum_{i=1}^{n}\log\frac{p_{x}(O_{i})}{p_{y}(O_{i})}.$
Taking expectation under $\mathcal{N}(x,\sigma^{2})^{\otimes L}$ gives $D_{\mathrm{KL}}(\mathcal{N}(x,\sigma^{2})^{\otimes L}\|\mathcal{N}(y,\sigma^{2})^{\otimes L})=L\,d(x,y)$. Substituting $x=\mu_{\mathrm{pers}}(w)$ and $y=\mu_{\mathrm{pers}}(w^{\prime})$ yields the Lemma. ∎
Proof of Lemma 5.2.
Let
$\widehat{\mu}_{L}(w)=\frac{1}{L}\sum_{i=1}^{L}O_{i},\quad\widehat{\mu}_{L}(w^{\prime})=\frac{1}{L}\sum_{i=1}^{L}O^{\prime}_{i},$
where
$O_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{\mathrm{pers}}(w),\sigma^{2})$
and
$O^{\prime}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{\mathrm{pers}}(w^{\prime}),\sigma^{2})$
are independent. Then $\widehat{\mu}_{L}(w)\sim\mathcal{N}(\mu_{\mathrm{pers}}(w),\sigma^{2}/L)$ and $\widehat{\mu}_{L}(w^{\prime})\sim\mathcal{N}(\mu_{\mathrm{pers}}(w^{\prime}),\sigma^{2}/L)$, hence
$\widehat{\Delta}:=\widehat{\mu}_{L}(w)-\widehat{\mu}_{L}(w^{\prime})\sim\mathcal{N}\!\left(\Delta(w,w^{\prime}),\,\frac{2\sigma^{2}}{L}\right).$
For any $t>0$, Markov’s inequality yields
$\mathbb{P}(\widehat{\Delta}\leq 0)$ $=\mathbb{P}\!\big(e^{-t\widehat{\Delta}}\geq 1\big)\leq\mathbb{E}[e^{-t\widehat{\Delta}}$]
$=\exp\!\left(-t\Delta(w,w^{\prime})+\frac{t^{2}}{2}\cdot\frac{2\sigma^{2}}{L}\right).$
Optimizing over $t$ gives $t^{\star}=\frac{L\Delta(w,w^{\prime})}{2\sigma^{2}}$, so
$\mathbb{P}(\widehat{\Delta}\leq 0)\leq\exp\!\left(-\frac{L\Delta(w,w^{\prime})^{2}}{4\sigma^{2}}\right).$
Thus it suffices that
$L\geq\frac{4\sigma^{2}}{\Delta(w,w^{\prime})^{2}}\log\frac{1}{\delta}$
Appendix B Extended discussions
B.1 AO and JPC - a counterexample
The equivalence in Lemma 4.3 highlights that (AO) is not merely technical: if micro-level information leaks (raw votes, rater identities, ordering, etc.), then two evaluation pipelines can induce the same aggregate kernel $Q(\cdot\mid w)$ yet still be distinguishable (and exploitable) by an adaptive method. The following minimal construction makes this necessity direction concrete.
Proposition B.1 (Violating (AO) can break JPC even when aggregate kernels match).
There exist two benchmarks $B$ and $B^{\prime}$ that induce the same reduced-form kernel $Q(\cdot\mid w)$ on the aggregate feedback space $\mathcal{O}$, but such that if the benchmark reveals the raw vote vector $(Z_{1},\dots,Z_{L})$ to the method (violating (AO)), then there is an adaptive method $\mathcal{A}$ whose aggregate transcript law $\mathbb{P}^{\mathcal{A}}(w_{1:T},o_{1:T})$ differs between $B$ and $B^{\prime}$. Consequently, the swap cannot be “just panel change” on the method interface as in Definition 2.
Proof.
We specify two benchmarks and an adaptive method.
Spaces and aggregation.
Let $\mathcal{W}=\{0,1\}$, $\mathcal{Z}=\{0,1\}$, $\mathcal{O}=\{0,1\}$, and take panel size $L=2$. Let the aggregation map be the XOR (disagreement) statistic
$\Gamma(z_{1},z_{2}):=z_{1}\oplus z_{2}\in\{0,1\}.$
Thus the aggregate output is $o=1$ iff the two individual votes disagree.
Two micro-instruments with identical aggregate kernels.
Let $p_{0}:=0.1$ and $p_{1}:=0.4$. Define benchmark $B$ so that, for each submitted artifact $w\in\{0,1\}$, the two micro-votes are independent Bernoulli draws
$Z_{1},Z_{2}\overset{i.i.d.}{\sim}\mathrm{Bernoulli}(p_{w}).$
Define benchmark $B^{\prime}$ identically except that each vote is flipped in distribution:
$Z^{\prime}_{1},Z^{\prime}_{2}\overset{i.i.d.}{\sim}\mathrm{Bernoulli}(1-p_{w}).$
(Panel sampling is irrelevant here; one can take a degenerate panel distribution and absorb everything into the micro-instrument.)
Now compute the induced reduced-form kernels on $\mathcal{O}$. Under $B$,
$\mathbb{P}_{B}(o=1\mid w)=\mathbb{P}_{B}(Z_{1}\neq Z_{2}\mid w)=2p_{w}(1-p_{w}).$
Under $B^{\prime}$,
$\mathbb{P}_{B^{\prime}}(o=1\mid w)=\mathbb{P}_{B^{\prime}}(Z^{\prime}_{1}\neq Z^{\prime}_{2}\mid w)=2(1-p_{w})p_{w}=2p_{w}(1-p_{w}).$
Hence the aggregate channels coincide exactly:
$Q_{B}(\cdot\mid w)=Q_{B^{\prime}}(\cdot\mid w)\qquad\forall\,w\in\mathcal{W}.$
Concretely, $Q(o=1\mid 0)=2(0.1)(0.9)=0.18$ and $Q(o=1\mid 1)=2(0.4)(0.6)=0.48$ for both benchmarks.
AO violation (raw-vote leakage) and an adaptive distinguisher.
Now suppose (AO) is violated and the benchmark releases the raw vote vector $(Z_{1},Z_{2})$ (or $(Z^{\prime}_{1},Z^{\prime}_{2})$) to the method in addition to the aggregate $o$. Consider the following horizon-$T=2$ adaptive method $\mathcal{A}$:
  • Round 1: submit $w_{1}=0$.
  • Observe the first raw vote and set $w_{2}:=Z_{1,1}$ (i.e., the round-1 vote of evaluator 1).
This is a valid adaptive strategy under the leaked interface, but it is not measurable with respect to the aggregate-only history $(w_{1},o_{1})$.
Under benchmark $B$, $\mathbb{P}(w_{2}=1)=\mathbb{P}(Z_{1,1}=1\mid w_{1}=0)=p_{0}=0.1$. Under benchmark $B^{\prime}$, $\mathbb{P}(w_{2}=1)=\mathbb{P}(Z^{\prime}_{1,1}=1\mid w_{1}=0)=1-p_{0}=0.9$. Therefore the distribution of the second-round submission $w_{2}$ differs between the two benchmarks even though the aggregate kernel $Q(\cdot\mid w)$ is identical.
Because $o_{2}\sim Q(\cdot\mid w_{2})$ in both benchmarks, this also induces a difference in the aggregate outcome at round 2:
$\mathbb{P}_{B}(o_{2}=1)$ $=\mathbb{P}_{B}(w_{2}=1)\,Q(o=1\mid 1)+\mathbb{P}_{B}(w_{2}=0)\,Q(o=1\mid 0)$
$=0.1\cdot 0.48+0.9\cdot 0.18=0.21,$
$\mathbb{P}_{B^{\prime}}(o_{2}=1)$ $=0.9\cdot 0.48+0.1\cdot 0.18=0.45.$
Thus the aggregate transcript laws $\mathbb{P}^{\mathcal{A}}_{B}(w_{1:2},o_{1:2})$ and $\mathbb{P}^{\mathcal{A}}_{B^{\prime}}(w_{1:2},o_{1:2})$ differ.
Conclusion.
If the swap $B\leftrightarrow B^{\prime}$ were JPC on the aggregate-only method interface in the sense of Definition 2, then (since $Q_{B}=Q_{B^{\prime}}$) the factorization would force every method’s aggregate transcript law to agree under $B$ and $B^{\prime}$. The above $\mathcal{A}$ contradicts this as soon as raw votes are released. Therefore leaking micro-level information (violating (AO)) can break JPC even when the aggregate kernels match. ∎
The same phenomenon occurs if the benchmark releases panel identities (or stable rater IDs): even if the aggregate score distribution $Q(\cdot\mid w)$ is unchanged, the extra identifier acts as a side channel that an adaptive method can condition on, producing different submission sequences and hence different aggregate transcripts. This is exactly why (AO) is a benchmark-hygiene requirement: it rules out side channels through which the method can distinguish evaluator implementations that are otherwise identical at the aggregate level.
B.2 Practical realism of method-blind evaluation (MB)
Assumption 2 is a protocol requirement: conditional on the submitted artifact $w$, the distribution of the returned aggregate feedback must not depend on who submitted $w$ or how it was produced. This is intentionally stronger than what one gets “by default” in many human-judgment settings, because there is extensive evidence that evaluations can shift when raters are exposed to provenance cues (e.g., explicit labels such as “AI-generated” vs. “human-written”) or other contextual metadata beyond the content being judged (Bogert et al., 2022; Osborne and Bailey, 2025; Levitt and List, 2011; de Quidt et al., 2018). In our framework, such provenance dependence is precisely an (MB) violation: two identical artifacts $w$ can induce different score distributions if the evaluator observes additional information correlated with the producing method. [^1] If provenance cues are literally part of the submitted artifact (e.g., the text contains “as an AI language model”), then any resulting penalty is not an (MB) violation, because it is a function of $w$ itself. The problematic case for (MB) is when the evaluator is shown extra metadata (model name, submitter identity, method label, timestamp-based context, etc.) that is not a function of $w$.
How realistic is (MB) in practice?
(MB) is best viewed as a design target that is often approximately achievable, but not automatic. It is most realistic in settings where the outcome is behavioral and passively recorded (classic online A/B tests), since users are typically not told which method produced what they see and outcomes like click-through or conversion are not direct subjective judgments (Kohavi et al., 2020). By contrast, in explicit rating / preference-judgment pipelines (crowd or expert), (MB) is fragile: even minimal provenance cues (labels, branding, prior beliefs about model quality, expectations about “humans vs. AI”) can create systematic shifts in ratings for the same artifact, violating (MB). This fragility is also consistent with the motivation for anonymous pairwise evaluation interfaces used in large-scale leaderboards (Chiang et al., 2024) and with documented ordering/labeling effects in LLM-judge-style protocols (Dominguez-Olmedo et al., 2024).
The key takeaway is that (MB) should not be read as an empirical claim about human invariance; it is a benchmark hygiene constraint that must be enforced (or at least audited) by the benchmark implementation.
B.3 Extension to heteroscedastic Gaussian reduced-form kernels
Lemma 5.1 assumes a homoscedastic Gaussian reduced-form kernel, i.e., $\mathrm{Var}(o\mid w)$ is constant across artifacts. In many benchmarks, however, the aggregate score variance depends on the submitted artifact: some artifacts elicit highly consistent micro-responses (low noise), while others are ambiguous or polarizing (high noise). This motivates the more general heteroscedastic Gaussian model
$Q_{\mathrm{pers}}(\cdot\mid w)=\mathcal{N}\!\big(\mu_{\mathrm{pers}}(w),\,\sigma_{\mathrm{pers}}^{2}(w)\big),\qquad\sigma_{\mathrm{pers}}(w)>0.$ (14)
(Here $\sigma_{\mathrm{pers}}^{2}(w)$ denotes the variance of the aggregate feedback draw $o\sim Q_{\mathrm{pers}}(\cdot\mid w)$. Any micro-level panel size has already been absorbed into this reduced-form variance.)
Lemma B.2 (KL divergence under heteroscedastic Gaussians).
Assume $\mathcal{O}=\mathbb{R}$ and (14). Then for any $w,w^{\prime}\in\mathcal{W}$,
$D_{\mathrm{KL}}\!\big(Q_{\mathrm{pers}}(\cdot\mid w)\,\big\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})\big)$
$=\frac{1}{2}\Bigg[\log\frac{\sigma_{\mathrm{pers}}^{2}(w^{\prime})}{\sigma_{\mathrm{pers}}^{2}(w)}+\frac{\sigma_{\mathrm{pers}}^{2}(w)}{\sigma_{\mathrm{pers}}^{2}(w^{\prime})}-1\Bigg]\;+\;\frac{\big(\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})\big)^{2}}{2\,\sigma_{\mathrm{pers}}^{2}(w^{\prime})}.$ (15)
Moreover, for $L$ i.i.d. aggregate evaluations (i.e., $Q_{\mathrm{pers}}(\cdot\mid w)^{\otimes L}$),
$D_{\mathrm{KL}}\!\Big(Q_{\mathrm{pers}}(\cdot\mid w)^{\otimes L}\,\Big\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})^{\otimes L}\Big)=L\,D_{\mathrm{KL}}\!\big(Q_{\mathrm{pers}}(\cdot\mid w)\,\big\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})\big).$
Equation (15) highlights two separable sources of distributional distinguishability: (i) a variance-mismatch term (the bracketed expression), which is zero iff $\sigma_{\mathrm{pers}}^{2}(w)=\sigma_{\mathrm{pers}}^{2}(w^{\prime})$, and (ii) a mean-separation term, which scales the squared mean gap by $1/\sigma_{\mathrm{pers}}^{2}(w^{\prime})$. In contrast to the homoscedastic case, the KL divergence is generally asymmetric: $D_{\mathrm{KL}}(Q(\cdot\mid w)\|Q(\cdot\mid w^{\prime}))\neq D_{\mathrm{KL}}(Q(\cdot\mid w^{\prime})\|Q(\cdot\mid w))$ when $\sigma_{\mathrm{pers}}^{2}(w)\neq\sigma_{\mathrm{pers}}^{2}(w^{\prime})$.
Proof of Lemma B.2.
Fix $w,w^{\prime}\in\mathcal{W}$ and abbreviate
$\mu:=\mu_{\mathrm{pers}}(w),\quad\mu^{\prime}:=\mu_{\mathrm{pers}}(w^{\prime}),\quad\sigma^{2}:=\sigma_{\mathrm{pers}}^{2}(w),\quad\tau^{2}:=\sigma_{\mathrm{pers}}^{2}(w^{\prime}).$
Let $P:=\mathcal{N}(\mu,\sigma^{2})$ and $Q:=\mathcal{N}(\mu^{\prime},\tau^{2})$ with Lebesgue densities
$p(o)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\!\left(-\frac{(o-\mu)^{2}}{2\sigma^{2}}\right),\qquad q(o)=\frac{1}{\sqrt{2\pi\tau^{2}}}\exp\!\left(-\frac{(o-\mu^{\prime})^{2}}{2\tau^{2}}\right).$
By definition,
$D_{\mathrm{KL}}(P\|Q)=\mathbb{E}_{O\sim P}\!\left[\log\frac{p(O)}{q(O)}\right].$
Compute the log-likelihood ratio:
$\log\frac{p(O)}{q(O)}$ $=\left(-\frac{1}{2}\log(2\pi\sigma^{2})-\frac{(O-\mu)^{2}}{2\sigma^{2}}\right)-\left(-\frac{1}{2}\log(2\pi\tau^{2})-\frac{(O-\mu^{\prime})^{2}}{2\tau^{2}}\right)$
$=\frac{1}{2}\log\frac{\tau^{2}}{\sigma^{2}}+\frac{(O-\mu^{\prime})^{2}}{2\tau^{2}}-\frac{(O-\mu)^{2}}{2\sigma^{2}}.$
Taking expectations under $O\sim\mathcal{N}(\mu,\sigma^{2})$, we use the identities
$\mathbb{E}[(O-\mu)^{2}]=\sigma^{2},\qquad\mathbb{E}[(O-\mu^{\prime})^{2}]=\mathrm{Var}(O)+(\mathbb{E}[O]-\mu^{\prime})^{2}=\sigma^{2}+(\mu-\mu^{\prime})^{2}.$
Hence
$D_{\mathrm{KL}}(P\|Q)$ $=\frac{1}{2}\log\frac{\tau^{2}}{\sigma^{2}}+\frac{1}{2\tau^{2}}\Big(\sigma^{2}+(\mu-\mu^{\prime})^{2}\Big)-\frac{1}{2\sigma^{2}}\sigma^{2}$
$=\frac{1}{2}\left[\log\frac{\tau^{2}}{\sigma^{2}}+\frac{\sigma^{2}}{\tau^{2}}-1\right]+\frac{(\mu-\mu^{\prime})^{2}}{2\tau^{2}}.$
Substituting back $\sigma^{2}=\sigma_{\mathrm{pers}}^{2}(w)$ and $\tau^{2}=\sigma_{\mathrm{pers}}^{2}(w^{\prime})$ gives (15).
For the product statement, note that $P^{\otimes L}$ and $Q^{\otimes L}$ have densities $p_{L}(o_{1:L})=\prod_{i=1}^{L}p(o_{i})$ and $q_{L}(o_{1:L})=\prod_{i=1}^{L}q(o_{i})$, so
$\log\frac{p_{L}(O_{1:L})}{q_{L}(O_{1:L})}=\sum_{i=1}^{L}\log\frac{p(O_{i})}{q(O_{i})}.$
Taking expectation under $O_{1:L}\sim P^{\otimes L}$ and using independence,
$D_{\mathrm{KL}}(P^{\otimes L}\|Q^{\otimes L})=\sum_{i=1}^{L}\mathbb{E}_{O_{i}\sim P}\!\left[\log\frac{p(O_{i})}{q(O_{i})}\right]=L\,D_{\mathrm{KL}}(P\|Q).$
This proves the lemma. ∎
Appendix C Experiment: Discriminability Calibration for the Persona-Based Ad Benchmark
This experiment provides a proof-of-concept discriminability audit for persona-simulation evaluation environment used in demonstrating the effectiveness of TextBO (Kang and Yoganarasimhan, 2025), a prompt optimization based self-improving AI method. Specifically, we focus on the ad optimization experimentation proposed in this paper.
In the ad optimization experiment, TextBO iteratively improves the prompt for ad image generation that is fed to an image model to produce the ad creative, using evaluation feedback to decide which prompt edits to keep. Concretely, it runs an automated prompt improvement loop: write a prompt $\rightarrow$ generate an ad image $\rightarrow$ evaluate it on a target audience $\rightarrow$ reflection $\rightarrow$ edit the prompt and repeat.
As in Kang and Yoganarasimhan (2025), we consider eight synthetic ad campaign scenarios that cover a diverse range of products across distinct categories, each defined by a creative brief that outlines the strategic and creative direction (see Web Appendix C.1 of Kang and Yoganarasimhan (2025) for the full creative briefs):
  • Scenario 1: “GreenBite,” a new plant-based burger patty.
  • Scenario 2: “AuraSonics X1,” high-end, noise-canceling wireless earbuds.
  • Scenario 3: “Odyssey E-SUV,” a new all-electric family SUV.
  • Scenario 4: “Oasis Eco-Lodge,” a secluded, luxury resort with beautiful natural surroundings.
  • Scenario 5: “Momentum,” a mobile-first banking app for freelancers and the gig economy.
  • Scenario 6: “MindGarden,” a subscription-based meditation and mindfulness app.
  • Scenario 7: “Aeterno,” a classic, automatic Swiss-made wristwatch with a heritage design.
  • Scenario 8: “SyncFlow,” a project management and collaboration software platform for remote teams.
For evaluation, TextBO utilizes Twin-2k-500 persona dataset (Toubia et al., 2025). For a fixed scenario and prompt, TextBO evaluates an ad image by: (i) sampling 200 personas from the Twin-2k-500 training split, (ii) conditioning a multimodal LLM judge on each persona’s survey answers, requesting an effectiveness rating on the 1–5 scale, (iii) converting the judge’s log-probabilities over {1,2,3,4,5} into a real-valued expected score per persona, and (iv) averaging over the 200 personas to produce a single scalar score. This single scalar is the only feedback used by the optimization method. We use Gemini 2.5 Flash with the same meta-prompt used in Kang and Yoganarasimhan (2025), which is provided in Figure 1.
Persona prompt for simulating ad effectiveness. SYSTEM: $\{$ You are an AI assistant. Your task is to answer the TASK as if you are the individual described in the ‘Persona Profile’ (which contains their past survey responses). Remain consistent with the persona’s past survey responses and stated characteristics. Carefully follow any instructions provided for the new question, including formatting requirements. } PERSONA DATA: $\{$
Which part of the United States do you currently live in?
Question Type: Single Choice
Options:
1 - Northeast (PA, NY, NJ, RI, CT, MA, VT, NH, ME)
2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
3 - South (TX, OK, AR, LA, KY, TN, MS, AL, WV, DC, MD, DE, VA, NC, SC, GA, FL)
4 - West (WA, OR, ID, MT, WY, CA, NV, UT, CO, AZ, NM)
5 - Pacific (HI, AK)
Answer: 2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
What is the highest level of schooling or degree that you have completed?
Question Type: Single Choice
Options:
1 - Less than high school
2 - High school graduate
3 - Some college, no degree
4 - Associate’s degree
5 - College graduate/some postgrad
6 - Postgraduate
Answer: 3 - Some college, no degree
$\ldots$ (Many other survey questions and answers) $\ldots$
Suppose you were given \$5 and had to offer to another (anonymous) person a way to split the money. The other person can either accept or reject your offer. If the other person accepts your offer, you would each receive the amount you proposed. If the other person rejects your offer, you would both receive \$0. How much would you offer to the other person?
Question Type: Single Choice
Options:
1 - \$0
2 - \$1
3 - \$2
4 - \$3
5 - \$4
6 - \$5
Answer: 3 - \$2
$\ldots$ (Many other survey questions and answers) $\ldots$
AD IMAGE: [image] TASK: Return only one item from [”1”,”2”,”3”,”4”,”5”] for ad effectiveness.
Effective Score Scale Definition:
1: Extremely Unlikely. The persona would actively ignore or be annoyed by this ad.
2: Unlikely. The persona would likely scroll past without a second thought.
3: Mediocre. It is hard to decide whether the personal would click or don’t click.
4: Likely. The persona is intrigued and has a good chance of clicking to learn more.
5: Extremely Likely. The persona is the ideal target; a click is almost certain.
No explanation. Just the score.
Figure 1: Meta-prompt for simulating the effectiveness of a given ad-persona combination.
Here, in this paper’s language, an artifact $w$ corresponds to [scenario + prompt + its generated ad image]. We denote an evaluation operation of an artifact $w$, which returns a single scalar $o\in\mathbb{R}$, as Eval(w). Repeating Eval(w) with fresh persona samples yields i.i.d. draws $o\sim Q_{\mathrm{pers}}(\cdot\mid w)$. As TextBO iteratively improves the prompt for ad image generation, we would like to answer the following question:
How many personas Eval(w) need to be confident that the improved prompt is better than the original prompt?
In this problem, $d_{W}(w,w^{\prime})$ naturally corresponds to the number of clause-level instruction edits needed to transform prompt $\pi$ into $\pi^{\prime}$. Therefore, we set the resolution $r=1$, i.e., we require the benchmark to reliably distinguish prompts that differ by at least one step of prompt improvement.
For each artifact $w\in W_{0}$, we run $n_{0}$ independent repeats of TextBO’s evaluation call:
$o_{1}(w),\dots,o_{n_{0}}(w)\;\;\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\;\;Q_{\mathrm{pers}}(\cdot\mid w),$
where each repeat resamples a fresh panel of $n_{0}=200$ personas (all other evaluation components are held fixed, including the judge prompt and model version). We compute the empirical mean and variance:
$\widehat{\mu}(w)=\frac{1}{n_{0}}\sum_{t=1}^{n_{0}}o_{t}(w),\qquad\widehat{\sigma}^{2}(w)=\frac{1}{n_{0}-1}\sum_{t=1}^{n_{0}}(o_{t}(w)-\widehat{\mu}(w))^{2}.$
For each scenario, for all pairs $w,w^{\prime}\in W_{0}$ with $d_{W}(w,w^{\prime})\geq r$, we compute the pairwise signal-to-noise proxy:
$\widehat{\mathrm{SNR}}(w,w^{\prime})\;=\;\frac{\big(\widehat{\mu}(w)-\widehat{\mu}(w^{\prime})\big)^{2}}{\widehat{\sigma}^{2}(w)+\widehat{\sigma}^{2}(w^{\prime})}.$
There are many possible combinations of $w,w^{\prime}$. Kang and Yoganarasimhan (2025) reported results for ten steps of prompt improvement from the initial prompt; we tested ten improvement variants for each step. This totals 100 TextBO prompt improvement cases per scenario. We then define a conservative discriminability estimate as a low quantile of these values, instead of taking a literal infimum:
$\widehat{\kappa}_{Q}\;=\;\mathrm{Quantile}_{q}\Big(\{\widehat{\mathrm{SNR}}(w,w^{\prime}):\;d_{W}(w,w^{\prime})\geq r\}\Big),$
with $q=0.05$. We report $\widehat{\kappa}_{Q}$ per scenario and use the minimum across scenarios as the overall conservative estimate. The predicted number of independent evaluation calls per prompt required to decide between two $r$-separated artifacts is
$n_{\mathrm{req}}\;=\;\left\lceil\frac{2}{\widehat{\kappa}_{Q}}\log\frac{1}{\delta}\right\rceil.$
Table 1 provides $\widehat{\kappa}_{Q}$ and $n_{\mathrm{req}}$ for each scenario. Across the scenarios, $n_{\text{req}}$ varies from 20 to 513, with an average of 204.38. This shows that (Kang and Yoganarasimhan, 2025)’s choice of random sampling 200 personas for training was a legitimate choice.
Table 1: $\widehat{\kappa}_{Q}$ and Sample Size for $\delta=0.05$
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Scenario 6 Scenario 7 Scenario 8 Average
$\widehat{\kappa}_{Q}$ 0.00508 0.02320 0.13178 0.00942 0.02955 0.13069 0.0046 0.07548 0.051225
$n_{\text{req}}$ 513 113 20 280 89 20 565 35 204.375